# FastSeg: Efficient Training-Free Open-Vocabulary Segmentation via Hierarchical Attention Refinement Method

### v1

@@ 4.2.1 Effects of different components
Table  3 presents an ablation study that evaluates the contribution of components in FastSeg. Removing the cross-attention and self-attention weighted aggregation modules is equivalent to using only a single-resolution attention map with the best performance (as shown in Table  5), without applying multi-resolution aggregation-specifically, $16\times 16$ for cross-attention and $32\times 32\times 32\times 32$ for self-attention. In contrast, removing the class prompt $\mathcal{P}^{\prime}$ means that cross-attention maps are computed solely based on the text prompt $\mathcal{P}$, without explicitly guiding the model to focus on candidate classes. In the original configuration, the class candidate list is appended to the image caption to form the complete prompt $\mathcal{P}$, allowing the model to utilize the appended candidate list to generate more targeted cross-attention maps. Among all components, self-attention refinement proves to be the most critical, as its removal leads to the most significant average performance drop (-15.1%), underscoring its importance in enhancing the spatial precision of class-aware attention. Both cross-attention and self-attention weighted aggregation also play important roles, as their removal results in noticeable performance degradation. On the other hand, class prompt design and test-time flipping (TTF) have relatively more minor effects-0.6% and 0.3%, respectively-yet consistently contribute to overall segmentation quality.
These findings validate the overall design of FastSeg, where each component serves a complementary role in improving segmentation performance while maintaining computational efficiency.
@@ 4.2.4 Effect of different attention scales
Table 5 presents a comprehensive study on the effect of different attention scales and fusion strategies in FastSeg. The results reveal several key findings. First, single-resolution attention maps, particularly those from the 16×16 resolution level, consistently yield higher segmentation performance across all datasets compared to other individual scales, suggesting that this resolution offers a favorable balance between semantic context and spatial detail. Second, when combining attention maps across multiple resolutions, both mean and weighted aggregation strategies improve overall performance, highlighting the complementary nature of multi-scale attention. Notably, weighted fusion outperforms mean fusion, as it enables the model to emphasize semantically rich and informative layers (e.g., 16×16 and 32×32) while down-weighting noisier or less discriminative ones (e.g., 64×64). The best overall performance—with an average mIoU of 64.0 / 36.2 / 31.3 on PASCAL VOC, Context, and COCO, respectively—is achieved using the weighted fusion of both cross- and self-attention maps, as indicated by the red-highlighted entry. Finally, attention maps at very low (8×8) or very high (64×64) resolutions individually underperform, reaffirming that mid-level resolutions are the most effective for capturing meaningful object-level features. These findings validate the design choice in FastSeg to apply multi-resolution attention fusion, with empirically optimized weights, to maximize segmentation accuracy across diverse datasets.
@@
To assess the impact of image description quality on FastSeg’s segmentation performance, we experiment with several text prompt generation strategies for constructing the reconstruction prompt $\mathcal{P}$. In addition to a simple text prompt—“A photo of {class candidates}”—which does not rely on any image captioning model, we evaluate a variety of automatic captioning methods. These include BLIP [47], used as the default captioning model in our framework, as well as more recent models such as BLIP-2 [48], GIT [49], ViT-GPT2 [50], and Gemini Flash 2.0 [51]. For Gemini, we consider two prompting strategies: Gemini(1) uses the instruction “Describe this image in one sentence” to generate a full description, whereas Gemini(2) uses “Describe the image in a short and generic sentence, avoid specific names or proper nouns” to generate more abstract, less detailed captions. As shown in Table 7, the model exhibits relatively stable performance across different captioning methods, indicating that FastSeg is not overly sensitive to variations in prompt content. Interestingly, we observe that overly detailed prompts (Gemini(1)) slightly degrade performance, whereas more simple prompts (Gemini(2)) yield more better results. Additionally, the simple text prompt also yields suboptimal performance, indicating that purely class-driven prompts without image-specific context are less effective in guiding segmentation. Overall, BLIP provides a balanced captioning quality that is neither too generic nor overly specific, making it a suitable choice for guiding attention in our FastSeg.
Table 7: Performance of FastSeg on the PASCAL VOC dataset using different methods for generating the text prompt $\mathcal{P}$.
@@ 5.1 Limitations
While FastSeg achieves strong segmentation performance and inference efficiency, it still faces several limitations. First, the method is sensitive to the quality of the candidate class list, which must be provided or predicted in advance; incomplete or noisy class sets can lead to false negatives or misclassifications. Second, FastSeg relies on static prompt encoding during inference—once the textual prompt is generated, it is not updated or refined adaptively, which may limit its flexibility in complex scenes. Third, as a training-free method, FastSeg’s performance is inherently constrained by the representational capacity of the pretrained diffusion backbone (e.g., Stable Diffusion); any limitations or biases in the underlying model directly affect segmentation quality and generalization.
In addition, we present several failure cases of FastSeg in Figure 10, which highlight the current limitations of our approach in challenging segmentation scenarios. The first three rows in the left column illustrate difficulties in accurately segmenting object shapes and boundaries, particularly in scenes containing multiple overlapping objects (e.g., the first two rows). These cases suggest limitations in spatial discrimination when semantic boundaries are ambiguous or tightly packed. The top row of the right column demonstrates FastSeg’s difficulty in detecting small objects, which may be overlooked due to their weak or missing representation in attention maps. Similarly, the second row of the right column shows segmentation errors when visually similar categories appear in close proximity, indicating challenges in category disambiguation under fine-grained visual similarity. Finally, the bottom row in both columns showcases cases where the quality of the candidate class list heavily influences segmentation performance. In these examples, the failure to predict or include correct class labels in the class prompt results in incomplete or incorrect segmentation, reaffirming FastSeg’s dependency on accurate open-set label generation.
Figure 10: Visualization of some failure cases by FastSeg
@@ 5.2 Conclusion
In this paper, we introduced FastSeg, a novel training-free open-vocabulary semantic segmentation (OVSS) method that leverages diffusion models for efficient and accurate segmentation. Our method combines (1+1)-step DDIM inversion, a dual-prompt mechanism, and a multi-resolution attention fusion strategy. Unlike prior approaches, FastSeg generates segmentation maps for all candidate classes in a single inference run, significantly reducing computational overhead. Additionally, we incorporate Test-Time Flipping (TTF) to improve prediction robustness. Extensive experiments on PASCAL VOC, PASCAL Context, and COCO Object benchmarks demonstrate that FastSeg achieves state-of-the-art performance among training-free diffusion-based methods, with the highest average mIoU and achive inference time under one seconds. These results highlight the practicality and scalability of FastSeg for real-world OVSS applications.
Building upon FastSeg, future research can explore several directions to further enhance segmentation quality and flexibility. One promising avenue is the development of adaptive candidate class generation mechanisms, which leverage contextual cues to dynamically refine class sets and reduce dependency on the auxiliary candidate class generation process. Additionally, extending FastSeg to support instance-level segmentation would enable more fine-grained visual understanding. With its high adaptability, accurate segmentation performance, and low segmentation latency, FastSeg provides a practical and deployable solution for automatic segmentation in real-world applications.

### v2

@@ 4.2.1 Effects of different components
Table  3 presents an ablation study that evaluates the contribution of components in SwiftSeg. Removing the cross-attention and self-attention weighted aggregation modules is equivalent to using only a single-resolution attention map with the best performance (as shown in Table  5), without applying multi-resolution aggregation-specifically, $16\times 16$ for cross-attention and $32\times 32\times 32\times 32$ for self-attention. In contrast, removing the class prompt $\mathcal{P}^{\prime}$ means that cross-attention maps are computed solely based on the text prompt $\mathcal{P}$, without explicitly guiding the model to focus on candidate classes. In the original configuration, the class candidate list is appended to the image caption to form the complete prompt $\mathcal{P}$, allowing the model to utilize the appended candidate list to generate more targeted cross-attention maps. Among all components, self-attention refinement proves to be the most critical, as its removal leads to the most significant average performance drop (-15.1%), underscoring its importance in enhancing the spatial precision of class-aware attention. Both cross-attention and self-attention weighted aggregation also play important roles, as their removal results in noticeable performance degradation. On the other hand, class prompt design and test-time flipping (TTF) have relatively more minor effects-0.6% and 0.3%, respectively-yet consistently contribute to overall segmentation quality.
These findings validate the overall design of SwiftSeg, where each component serves a complementary role in improving segmentation performance while maintaining computational efficiency.
@@ 4.2.4 Effect of different attention scales
Table 5 presents a comprehensive study on the effect of different attention scales and fusion strategies in SwiftSeg. The results reveal several key findings. First, single-resolution attention maps, particularly those from the 16×16 resolution level, consistently yield higher segmentation performance across all datasets compared to other individual scales, suggesting that this resolution offers a favorable balance between semantic context and spatial detail. Second, when combining attention maps across multiple resolutions, both mean and weighted aggregation strategies improve overall performance, highlighting the complementary nature of multi-scale attention. Notably, weighted fusion outperforms mean fusion, as it enables the model to emphasize semantically rich and informative layers (e.g., 16×16 and 32×32) while down-weighting noisier or less discriminative ones (e.g., 64×64). The best overall performance—with an average mIoU of 64.0 / 36.2 / 31.3 on PASCAL VOC, Context, and COCO, respectively—is achieved using the weighted fusion of both cross- and self-attention maps, as indicated by the red-highlighted entry. Finally, attention maps at very low (8×8) or very high (64×64) resolutions individually underperform, reaffirming that mid-level resolutions are the most effective for capturing meaningful object-level features. These findings validate the design choice in SwiftSeg to apply multi-resolution attention fusion, with empirically optimized weights, to maximize segmentation accuracy across diverse datasets.
@@
To assess the impact of image description quality on SwiftSeg’s segmentation performance, we experiment with several text prompt generation strategies for constructing the reconstruction prompt $\mathcal{P}$. In addition to a simple text prompt—“A photo of {class candidates}”—which does not rely on any image captioning model, we evaluate a variety of automatic captioning methods. These include BLIP [47], used as the default captioning model in our framework, as well as more recent models such as BLIP-2 [48], GIT [49], ViT-GPT2 [50], and Gemini Flash 2.0 [51]. For Gemini, we consider two prompting strategies: Gemini(1) uses the instruction “Describe this image in one sentence” to generate a full description, whereas Gemini(2) uses “Describe the image in a short and generic sentence, avoid specific names or proper nouns” to generate more abstract, less detailed captions. As shown in Table 7, the model exhibits relatively stable performance across different captioning methods, indicating that SwiftSeg is not overly sensitive to variations in prompt content. Interestingly, we observe that overly detailed prompts (Gemini(1)) slightly degrade performance, whereas more simple prompts (Gemini(2)) yield more better results. Additionally, the simple text prompt also yields suboptimal performance, indicating that purely class-driven prompts without image-specific context are less effective in guiding segmentation. Overall, BLIP provides a balanced captioning quality that is neither too generic nor overly specific, making it a suitable choice for guiding attention in our SwiftSeg.
Table 7: Performance of SwiftSeg on the PASCAL VOC dataset using different methods for generating the text prompt $\mathcal{P}$.
@@ 5.1 Limitations
While SwiftSeg achieves strong segmentation performance and inference efficiency, it still faces several limitations. First, the method is sensitive to the quality of the candidate class list, which must be provided or predicted in advance; incomplete or noisy class sets can lead to false negatives or misclassifications. Second, SwiftSeg relies on static prompt encoding during inference—once the textual prompt is generated, it is not updated or refined adaptively, which may limit its flexibility in complex scenes. Third, as a training-free method, SwiftSeg’s performance is inherently constrained by the representational capacity of the pretrained diffusion backbone (e.g., Stable Diffusion); any limitations or biases in the underlying model directly affect segmentation quality and generalization.
In addition, we present several failure cases of SwiftSeg in Figure 10, which highlight the current limitations of our approach in challenging segmentation scenarios. The first three rows in the left column illustrate difficulties in accurately segmenting object shapes and boundaries, particularly in scenes containing multiple overlapping objects (e.g., the first two rows). These cases suggest limitations in spatial discrimination when semantic boundaries are ambiguous or tightly packed. The top row of the right column demonstrates SwiftSeg’s difficulty in detecting small objects, which may be overlooked due to their weak or missing representation in attention maps. Similarly, the second row of the right column shows segmentation errors when visually similar categories appear in close proximity, indicating challenges in category disambiguation under fine-grained visual similarity. Finally, the bottom row in both columns showcases cases where the quality of the candidate class list heavily influences segmentation performance. In these examples, the failure to predict or include correct class labels in the class prompt results in incomplete or incorrect segmentation, reaffirming SwiftSeg’s dependency on accurate open-set label generation.
Figure 10: Visualization of some failure cases by SwiftSeg
@@ 5.2 Conclusion
In this paper, we introduced SwiftSeg, a novel training-free open-vocabulary semantic segmentation (OVSS) method that leverages diffusion models for efficient and accurate segmentation. Our method combines (1+1)-step DDIM inversion, a dual-prompt mechanism, and a multi-resolution attention fusion strategy. Unlike prior approaches, SwiftSeg generates segmentation maps for all candidate classes in a single inference run, significantly reducing computational overhead. Additionally, we incorporate Test-Time Flipping (TTF) to improve prediction robustness. Extensive experiments on PASCAL VOC, PASCAL Context, and COCO Object benchmarks demonstrate that SwiftSeg achieves state-of-the-art performance among training-free diffusion-based methods, with the highest average mIoU and achive inference time under one seconds. These results highlight the practicality and scalability of SwiftSeg for real-world OVSS applications.
Building upon SwiftSeg, future research can explore several directions to further enhance segmentation quality and flexibility. One promising avenue is the development of adaptive candidate class generation mechanisms, which leverage contextual cues to dynamically refine class sets and reduce dependency on the auxiliary candidate class generation process. Additionally, extending SwiftSeg to support instance-level segmentation would enable more fine-grained visual understanding. With its high adaptability, accurate segmentation performance, and low segmentation latency, SwiftSeg provides a practical and deployable solution for automatic segmentation in real-world applications.